OpenStudy (anonymous):

simplify (y-8)/(y-9) - (y+1)/(y+9) + (y-27)/(y^2-81) please put the steps, i am not quite sure where to start on this one, thanks!

OpenStudy (saifoo.khan):

Hello.

OpenStudy (anonymous):

hi saifoo, i got (2y+12)/(y+9), is this right?

OpenStudy (saifoo.khan):

you are almost correct.

OpenStudy (saifoo.khan):

10/(y+9)

OpenStudy (anonymous):

the first fraction needs to be mutliplied by x-9 over itself to get the LCD; the second fraction must be mutliplird by y+9 over itself to get the LCD; the third fractino already has the LCD;then I combined fractions after FOIL-ing\[\frac{y^2+y-72-y^2+8y+9+y-27}{(y-9)(y+9)}\] (i will continue on next post)

OpenStudy (saifoo.khan):

Mondolio when you are free, please have a look.

OpenStudy (saifoo.khan):

i mean please have a look in my question

OpenStudy (anonymous):

and mandolino, could you please check my other one i have posted about answering in radical form?

OpenStudy (anonymous):

btw, the 99 above is supposed to be a 90 \[=\frac{10y-90}{(x-9)(x+9)}\]factor out a 10 in the numerator\[=\frac{10\cancel{(y-9)}}{\cancel{(x-9)}(x+9)}=\frac{10}{x+9}\]which is wha saifoo got about 10 minutes ago!

OpenStudy (anonymous):

ahhhh ok, thank you

OpenStudy (anonymous):

correction: two posts ago the +9 is supposed to be a -9 (forget about the 99 and 90 stuff above!)

OpenStudy (anonymous):

i'll go look now at the radcial form one